How much would it cost to carpet a room that is 25 feet by 28 feet if carpeting costs $23.75 per square yard?
Answer by rfer(16322)

25*28=700 sq ft
1 yd = 9 sq ft
700/9=77.8 yds
77.8*23.75=$1847.75